OpenStudy (mathhelp346):

A rectangular garden is designed so that its length is twice its width. The number of feet in the perimeter of the garden is equal to the number of square feet in it area. What are the dimensions of the garden?

OpenStudy (anonymous):

Let the width of the rectangle be x. The length will be 2x, since it is given that the length is twice its width. Then we also know that the perimeter of the garden is equal to its area. Perimeter = area 2 length + 2 width = length x width Substituting in the x value, 2(2x) + 2x = 2x * x 4x+2x = 2x^2 6x = 2x^2 6= 2x x= 3 Dimensions of the garen, width= 3feet length = 6 feet

OpenStudy (mathhelp346):

so the perimeter=area is the equation? why?

OpenStudy (mathhelp346):

or is it just because it makes sense? ._.